# Enhancing nanocrystal superlattice self-assembly near a metastable liquid binodal

### v2

@@ 0.0.1 Keywords:
We further specify the interaction ranges and well depths by comparing the self-assembly of PbS NCs in this work with that of previously studied electrostatically stabilized Au NCs coated with the same thiostannate surface ligands. When Au NCs are quenched in the same way as in this work, they form SLs much more rapidly (within seconds), the SLs are more disordered, and there is no evidence from in situ measurements of a metastable liquid phase22, 45. The faster kinetics imply that Au NC interparticle potentials have larger values of $u_{0}$ than PbS NC interparticle potentials for similar quench conditions despite also being short-range. To explain the different kinetics using findings in Ref.3, we propose that the PbS NC interparticle potentials likely have $u_{0}$$\sim$2-3 $k_{\mathrm{B}}T$ in order to generate highly ordered SLs, while the Au NC interparticle potentials likely have $u_{0}$$\sim$3-6 $k_{\mathrm{B}}T$. Such a difference would imply that, for the same change in external control parameter, such as solution ionic strength, the change in Au NC $u_{0}$ is larger than the change in PbS NC $u_{0}$. This difference in sensitivity is likely caused by PbS NCs having weaker vdW attraction than Au NCs as well as the fact that solvents with a more strongly screening, higher dielectric constant were used for the PbS NC system compared to the Au NC system.22 The presence of a metastable liquid phase in the PbS NC SL self-assembly and the absence of observed liquid in the Au NC case also suggests that, in combination with qualitative predictions from Derjaguin-Landau-Verwey-Overbeek (DLVO) theory46, Au NC interactions are even shorter in range with respect to the effective size of an Au NC than PbS NC interactions are with respect to their effective size (Figure S11). This possibility could be due to Au NCs having larger effective sizes relative to the bare NC core than PbS NCs have due to their greater surface ligand coverage, which is expected due to their higher NC-to-solvent dielectric constant ratio.47 In summary, while both systems have accessible short-range interaction regimes whose attractive wells are not so deep nor so short15 as to reach commonly-observed gelation,40, 48, 41 we infer that PbS NC interactions are softer and longer in range than Au NC interactions when employing the same electrostatically stabilizing ligands and salts. The sensitivity of the PbS interparticle potentials to external control parameters ultimately enables the fine control necessary to tune the thermodynamic state, self-assembly pathway, and kinetics in a nanoscale system.
@@
In Figure 3c, the liquid and SL appear concurrently within experimental error, which is not formally expected from the one-step or the two-step pathways. This observation motivates an Avrami analysis on a series of $\sim$10 experiments at a constant quench depth of $\lambda$ = 0.296 nm (1.2 M) and varied volume fraction (Figure 3e orange, Figure S12) to describe SL nucleation and growth rates as a function of $\phi_{0}$ in this region of the parameter space (yellow in Figure 4a). Interestingly, repeating this same procedure for the rate of emergence of the liquid phase vs $\phi_{0}$, despite the more significant challenge of obtaining the rate constants for the more smoothly varying structure factor of the liquid, especially at low $\phi_{0}$ where the amplitude of this contribution to $I(q)$ is quite small, reproduces an essentially similar trend (blue). To combine the data from different volume fractions to obtain a much more robust assessment of the growth rates than can be obtained from a single Avrami fit on its own, we fit the Avrami rate constant, $k$ vs $\phi_{0}$ for the liquid and SL phases to a function arising from classical nucleation theory (dashed curves, see Methods). Given these trends of Avrami rate constant $k$ vs. $\phi_{0}$ for the liquid and SL phases match reasonably well over several orders of magnitude (Figure 4a), we propose that kinetics in this region are one-step like in that they are limited by a transition from a colloidal to a dense phase, irrespective of its order. The colloid-to-liquid transition is rate-limiting, and the liquid to SL transition is fast, suggesting a free energy landscape such as that sketched at right in Figure 3f.1 We corroborated this finding with in situ dynamic light scattering (DLS) measurements during self-assembly, which show that the average size of individual dense clusters grows as $\sim t^{1/2}$, consistent with expectations from nucleation and growth kinetics49 (see Figure S13 and associated text). These kinetics are also consistent with Wedekind et al.’s simulations of crystal nucleation close to a colloid-liquid phase transition7. They found that crystal (SL) formation in the region between the colloid-metastable liquid binodal and its spinodal curve appears to follow one-step kinetics despite the necessary presence of the liquid state; they regard this regime as exhibiting “effective one-step kinetics.” Alternatively, one might consider the possibility that one- and two-step pathways co-occur with similar rates, i.e., compete with one another, in this region of the phase diagram. The “effective one-step” mechanism is, however, more likely, not only because of correspondence with this simulation work, but also because we observe that the two-step pathway dominates kinetically deeper beneath the binodal (Figure 3d) and because it seems less plausible for independent one- and two-step pathways to have essentially identical rates over the entire range studied with the Avrami analysis. Ultimately, Figure 3e layers kinetic information on top of thermodynamic information to succinctly summarize the variety of behaviors accessible with this system, and can be used to prescribe conditions for diverse self-assembly outcomes.
@@
Through this establishment of possible kinetic pathways, we find that the liquid phase provides new opportunities to optimize self-assembly. First, liquid appears before or in tandem with SL for all quenches beneath the colloid-liquid binodal in Figure 3e, suggesting that the two-step and effective one-step processes are faster than the one-step process in this region, consistent with findings from simulations1, 7, 3 and protein crystallization12, 13 experiments (Figure 4b). It is still possible for the rate of the one-step process outside of the colloid-metastable coexistence liquid region to be faster than the rate of the effective one-step or two-step processes inside the colloid-metastable liquid coexistence region if larger NC volume fractions are used in the one-step process. Increasing the NC volume fraction to increase the rate of a one-step process, however, often introduces disorder (purple curve, Figure 4c)3, 22.
@@
Key factors for increasing crystallinity concomittantly with crystallization driving force. Combining the above insights about the interaction potential range and depth with the effective one-step kinetic pathway uncovered to achieve high rates of formation concurrently with high SL quality provides a prescription for the ultimate control of SL self-assembly, which should be translatable to other nanoscale systems. For short-range-interacting systems, the kinetic parameter space associated with the effective one-step region is determined by the location of the colloid-metastable liquid binodal, which can be controlled by the range of the interparticle interactions. Achieving short-range interactions via control of electrostatics is a highly desirable design strategy: tuning electrostatic repulsion relative to empirically determined van der Waals attraction provides access to a broad, rich parameter space in which to tune kinetic rates. It furthermore allows the yield, size, and crystallinity of self-assembly products to be controlled. Additionally, as illustrated by comparing the Au and PbS NC SL assembly, even within the context of electrostatic stabilization and short-range interacting particles, accessing the effective one-step regime to simultaneously optimize crystallization rate and product quality requires the ability to carefully tune the short-range potential. We propose that the NC-to-solvent dielectric constant ratio, which determines the effective NC size and associated interaction potential range, is critical to tune. Specifically, it must be sufficiently high to achieve short-range interactions that suppress the colloid-liquid binodal beneath the colloid-SL binodal while the solvent dielectric constant is kept high enough that, upon quenching via an increase of the ionic strength, the potential depth becomes no greater than $\sim$3 $k_{\mathrm{B}}T$ and the range does not become too small, so as to avoid gelation15. This prescription is achieved with the PbS NC system using a dielectric ratio of $\sim$2, as compared to the Au system whose ratio is unbounded. A range of ratios are accessible by varying the relative proportions of miscible solvents, such as the NMF and DMF used here, so that the accessible quench potential depths that are fine-tuned by the ionic strength surround the identified few-$k_{\mathrm{B}}T$ window. Examples are shown in Figure S16. These design principles can be applied directly to other colloidal and protein13, 9 systems to test their generality and optimize their utility in promoting intricate material design on the nanoscale and also at the microscale, provided index matching is no longer necessary,15 for example by having a scattering technique take the place of optical microscopy. It furthermore suggests the possibility of using inorganic nanoscale systems as models for liquid-liquid phase separation of biomolecules50 when X-ray probes can provide additional, higher-resolution, insight that would be incompatible with proteins themselves. Furthermore, using similar in situ measurement and analysis, predictive optimal protocols for assembly and function of a wide range of other complex systems ranging from materials for catalysis51 to drug delivery52 to energy storage53 will also come into reach.
@@ Methods
In situ SAXS experiments. SAXS data were collected at the Stanford Synchrotron Radiation Lightsource (SSRL) at beamline 1-5 with a photon energy of 15 keV and beam size of 600 $\times$ 600 $\mu$m and at the Advanced Light Source at beamline 7.3.3 with a photon energy of 10 keV and beam size of $\sim$500 $\mu$m (see Figure S2). Stock solutions of 5.8$\pm$0.3 nm PbS NCs with $\mathrm{K}_{4}\mathrm{Sn}_{2}\mathrm{S}_{6}$ ligands (see Supporting Information) in N-methylformamide (NMF) and 0.5 M $\mathrm{K}_{3}\mathrm{AsS}_{4}$ salt in NMF were prepared following a procedure previously outlined17. In a nitrogen filled glovebox, 400-500 $\mu$Ls of a solution of 5.8 nm PbS NCs with $\mathrm{K}_{4}\mathrm{Sn}_{2}\mathrm{S}_{6}$ ligands in mixtures of NMF and N,N-dimethylformamide (DMF) was loaded into a 2 mm path length quartz cuvette with custom 200 $\mu$m thick windows. A small stir bar was placed into the cuvette in the plane of the cuvette and the cuvette was then sealed using a rubber septum and parafilm. A syringe preloaded with a solution of 0.5 M $\mathrm{K}_{3}\mathrm{AsS}_{4}$ in NMF was attached to the cuvette via Teflon tubing through the septum. The tubing-septum interface was sealed with epoxy. The gas-tight apparatus was carefully moved into the beam path and the syringe placed onto a New Era syringe pump (model NE-1000). X-ray scattering data were collected continuously while stirring the solution using a magnetic stirrer from Ultrafast Systems. All X-ray scattering patterns were collected using $\sim$0.1--1 s exposures at a rate of one pattern every $\sim$0.1--5 s. For each in situ experiment, after about 5 min of data acquisition, the excess salt in NMF solution was injected using the syringe pump at a rate of $\sim$1 mL/s. The injection took $\sim$ 1 - 10 s depending on how much salt was added. The ionic strength of the solution post-quench is $I=\frac{1}{2}\sum_{n=1}^{N}c_{n}z_{n}^{2}$, where $c_{n}$ is the concentration of ion species $n$ in molar, $z_{n}$ is the valency of ion $n$, and $N$ is the number of different ion species in solution. The Debye length of the solution post-quench is $\lambda=\sqrt{\frac{\epsilon_{r}\epsilon_{0}k_{\mathrm{B}}T}{\sum_{n=1}^{N}c_{n}z_{n}^{2}e^{2}}}$ where $\epsilon_{r}$ is the solvent dielectric constant of 88, $\epsilon_{0}$ is the vacuum permittivity, and $e$ is the elemental charge. The apparatus was kept at room temperature (see Supporting Information for additional temperature considerations). Data were continuously acquired after injection for up to two hours. SAXS patterns of cuvettes filled with NMF/DMF and varying amounts of $\mathrm{K}_{3}\mathrm{AsS}_{4}$ salt were taken for background subtraction (Figure S5).
@@
Modeling of SAXS patterns. We model the background-subtracted scattered intensity as $I(q)=I_{\mathrm{colloid}}(q)+I_{\mathrm{liquid}}(q)+I_{\mathrm{SL}}(q)+I_{\mathrm{LQ}}(q)$, where $I_{\mathrm{colloid}}(q)$ is the scattered intensity from dilute, polydisperse, hard spheres, $I_{\mathrm{liquid}}(q)$ is the scattered intensity at high $q$ from a liquid of hard spheres, $I_{\mathrm{SL}}(q)$ is the scattered intensity at high $q$ from fcc SLs, and $I_{\mathrm{LQ}}(q)$ is the scattering at low $q$ due to the finite size of liquid and SL clusters. For $I_{\mathrm{colloid}}(q)$, $I_{\mathrm{SL}}(q)$, and $I_{\mathrm{LQ}}(q)$, we use models previously developed22. To compute $I_{\mathrm{liquid}}(q)$, we solve the Ornstein-Zernicke equation using the Percus-Yevick closure relation for the case of hard spheres54, 55. To explicitly compute $I_{\mathrm{colloid}}(q)$ and $I_{\mathrm{liquid}}(q)$ we use xrsdkit (https://github.com/scattering-central/xrsdkit). See supporting information for further details on the components of the model.
@@
Avrami analysis. We analyzed kinetics of liquid and SL formation using the Avrami equation, 57 $V(t)=1-\exp(kt^{n})$, where $V(t)$ is the volume fraction of the new phase. The Avrami exponent, $n$, which is typically between 1 and 4, describes the dimensionality of the system, with a contribution of 1 for nucleation and a contribution of 1 for each of the growth dimensions. We assume homogeneous nucleation and growth in 3 dimensions and thus expect an Avrami exponent of $n\sim 4$. For $n=4$, the Avrami rate constant, $k$, is $k=\frac{1}{3}k_{n}k_{g}^{3}$, where $k_{n}$ is the nucleation rate and $k_{g}$ is the growth rate in a single dimension. We use the fractions of NCs in the liquid and SL phases obtained from the scattering invariant calculation described above as proxies for $V(t)$, and fit these to the Avrami equation to extract $k$. Including density dependencies in the nucleating barrier and particle collision frequencies, classical nucleation theory predicts that $k$ depends on $\phi_{0}$ as $k=c\phi_{0}^{3}\exp(-\frac{a}{(\ln(\phi_{0}/\phi_{\mathrm{colloid}}))^{2}})$. 58, 59, 60 The parameters $a$, which depends on temperature and surface energy of the interface of the nucleus and the colloidal phase, $c$, which arises from the growth rate and the kinetic prefactor to the nucleation rate, and $\phi_{\mathrm{colloid}}$, which is the volume fraction of the colloidal phase at equilibrium, do not vary with $\phi_{0}$. See supporting information for further details.

### v4

@@ 0.0.1 Keywords:
We further specify the interaction ranges and well depths by comparing the self-assembly of PbS NCs in this work with that of previously studied electrostatically stabilized Au NCs coated with the same thiostannate surface ligands. When Au NCs are quenched in the same way as in this work, they form SLs much more rapidly (within seconds), the SLs are more disordered, and there is no evidence from in situ measurements of a metastable liquid phase23, 46. The faster kinetics imply that Au NC interparticle potentials have larger values of $u_{0}$ than PbS NC interparticle potentials for similar quench conditions despite also being short-range. To explain the different kinetics using findings in Ref.3, we propose that the PbS NC interparticle potentials likely have $u_{0}$$\sim$2-3 $k_{\mathrm{B}}T$ in order to generate highly ordered SLs, while the Au NC interparticle potentials likely have $u_{0}$$\sim$3-6 $k_{\mathrm{B}}T$. Such a difference would imply that, for the same change in external control parameter, such as solution ionic strength, the change in Au NC $u_{0}$ is larger than the change in PbS NC $u_{0}$. This difference in sensitivity is likely caused by PbS NCs having weaker vdW attraction than Au NCs as well as the fact that solvents with a more strongly screening, higher dielectric constant were used for the PbS NC system compared to the Au NC system.23 The presence of a metastable liquid phase in the PbS NC SL self-assembly and the absence of observed liquid in the Au NC case also suggests that, in combination with qualitative predictions from Derjaguin-Landau-Verwey-Overbeek (DLVO) theory47, Au NC interactions are even shorter in range with respect to the effective size of an Au NC than PbS NC interactions are with respect to their effective size (Figure SLABEL:SI-Au). This possibility could be due to Au NCs having larger effective sizes relative to the bare NC core than PbS NCs have due to their greater surface ligand coverage, which is expected due to their higher NC-to-solvent dielectric constant ratio.48 In summary, while both systems have accessible short-range interaction regimes whose attractive wells are not so deep nor so short15 as to reach commonly-observed gelation,41, 49, 42 we infer that PbS NC interactions are softer and longer in range than Au NC interactions when employing the same electrostatically stabilizing ligands and salts. The sensitivity of the PbS interparticle potentials to external control parameters ultimately enables the fine control necessary to tune the thermodynamic state, self-assembly pathway, and kinetics in a nanoscale system.
@@
In Figure 3c, the liquid and SL appear concurrently within experimental error, which is not formally expected from the one-step or the two-step pathways. This observation motivates an Avrami analysis on a series of $\sim$10 experiments at a constant quench depth of $\lambda$ = 0.296 nm (1.2 M) and varied volume fraction (Figure 3e orange, Figure SLABEL:SI-Avrami) to describe SL nucleation and growth rates as a function of $\phi_{0}$ in this region of the parameter space (yellow in Figure 4a). Interestingly, repeating this same procedure for the rate of emergence of the liquid phase vs $\phi_{0}$ reproduces an essentially similar trend (blue). This similarity is observed despite the more significant challenge of obtaining the rate constants for the more smoothly varying structure factor of the liquid, especially at low $\phi_{0}$ where the amplitude of this contribution to $I(q)$ is quite small. To combine the data from different volume fractions to obtain a much more robust assessment of the growth rates than can be obtained from a single Avrami fit on its own, we fit the Avrami rate constant, $k$ vs $\phi_{0}$ for the liquid and SL phases to a function arising from classical nucleation theory (dashed curves, see Methods). Given these trends of Avrami rate constant $k$ vs. $\phi_{0}$ for the liquid and SL phases match reasonably well over several orders of magnitude (Figure 4a), we propose that kinetics in this region are one-step like in that they are limited by a transition from a colloidal to a dense phase, irrespective of its order. The colloid-to-liquid transition is rate-limiting, and the liquid to SL transition is fast, suggesting a free energy landscape such as that sketched at right in Figure 3f.1 We corroborated this finding with in situ dynamic light scattering (DLS) measurements during self-assembly, which show that the average size of individual dense clusters grows as $\sim t^{1/2}$, consistent with expectations from nucleation and growth kinetics50 (see Figure SLABEL:SI-DLS and associated text). These kinetics are also consistent with Wedekind et al.’s simulations of crystal nucleation close to a colloid-liquid phase transition7. They found that crystal (SL) formation in the region between the colloid-metastable liquid binodal and its spinodal curve appears to follow one-step kinetics despite the necessary presence of the liquid state; they regard this regime as exhibiting “effective one-step kinetics.” Alternatively, one might consider the possibility that one- and two-step pathways co-occur with similar rates, i.e., compete with one another, in this region of the phase diagram. The “effective one-step” mechanism is, however, more likely, not only because of correspondence with this simulation work. First, we observe that the two-step pathway dominates kinetically deeper beneath the binodal (Figure 3d). Second, it seems less plausible for independent one- and two-step pathways to have essentially identical rates over the entire range studied with the Avrami analysis. Ultimately, Figure 3e layers kinetic information on top of thermodynamic information to succinctly summarize the variety of behaviors accessible with this system, and can be used to prescribe conditions for diverse self-assembly outcomes.
@@
Through this establishment of possible kinetic pathways, we find that the liquid phase provides new opportunities to optimize self-assembly. First, liquid appears before or in tandem with SL for all quenches beneath the colloid-liquid binodal in Figure 3e, suggesting that the two-step and effective one-step processes are faster than the one-step process in this region, consistent with findings from simulations1, 7, 3 and protein crystallization12, 13 experiments (Figure 4b). It is still possible for the rate of the one-step process outside of the colloid-metastable coexistence liquid region to be faster than the rate of the effective one-step or two-step processes inside the colloid-metastable liquid coexistence region if larger NC volume fractions are used in the one-step process. Increasing the NC volume fraction to increase the rate of a one-step process, however, often introduces disorder (purple curve, Figure 4c)3, 23.
@@
Key factors for increasing crystallinity concomittantly with crystallization driving force. Combining the above insights about the interaction potential range and depth with the effective one-step kinetic pathway uncovered to achieve high rates of formation concurrently with high SL quality provides a prescription for the ultimate control of SL self-assembly, which should be translatable to other nanoscale systems. For short-range-interacting systems, the kinetic parameter space associated with the effective one-step region is determined by the location of the colloid-metastable liquid binodal, which can be controlled by the range of the interparticle interactions. Achieving short-range interactions via control of electrostatics is a highly desirable design strategy: tuning electrostatic repulsion relative to empirically determined van der Waals attraction provides access to a broad, rich parameter space in which to tune kinetic rates. It furthermore allows the yield, size, and crystallinity of self-assembly products to be controlled. Additionally, as illustrated by comparing the Au and PbS NC SL assembly, even within the context of electrostatic stabilization and short-range interacting particles, accessing the effective one-step regime to simultaneously optimize crystallization rate and product quality requires the ability to carefully tune the short-range potential. We propose that the NC-to-solvent dielectric constant ratio, which determines the effective NC size and associated interaction potential range, is critical to tune. Specifically, it must be sufficiently high to achieve short-range interactions that suppress the colloid-liquid binodal beneath the colloid-SL binodal while the solvent dielectric constant is kept high enough that, upon quenching via an increase of the ionic strength, the potential depth becomes no greater than $\sim$3 $k_{\mathrm{B}}T$ and the range does not become too small, so as to avoid gelation15. This prescription is achieved with the PbS NC system using a dielectric ratio of $\sim$2, as compared to the Au system whose ratio is unbounded. A range of ratios are accessible by varying the relative proportions of miscible solvents, such as the NMF and DMF used here, so that the accessible quench potential depths that are fine-tuned by the ionic strength surround the identified few-$k_{\mathrm{B}}T$ window. Examples are shown in Figure SLABEL:SI-dielectric. These design principles can be applied directly to other colloidal and protein13, 9 systems to test their generality and optimize their utility in promoting intricate material design on the nanoscale and also at the microscale, provided index matching is no longer necessary,15 for example by having a scattering technique take the place of optical microscopy. It furthermore suggests the possibility of using inorganic nanoscale systems as models for liquid-liquid phase separation of biomolecules51 when X-ray probes can provide additional, higher-resolution, insight that would be incompatible with proteins themselves. Furthermore, using similar in situ measurement and analysis, predictive optimal protocols for assembly and function of a wide range of other complex systems ranging from materials for catalysis52 to drug delivery53 to energy storage54 will also come into reach.
@@ Methods
In situ SAXS experiments. SAXS data were collected at the Stanford Synchrotron Radiation Lightsource (SSRL) at beamline 1-5 with a photon energy of 15 keV and beam size of 600 $\times$ 600 $\mu$m and at the Advanced Light Source at beamline 7.3.3 with a photon energy of 10 keV and beam size of $\sim$500 $\mu$m (see Figure SLABEL:SI-aparatus). Stock solutions of 5.8$\pm$0.3 nm PbS NCs with $\mathrm{K}_{4}\mathrm{Sn}_{2}\mathrm{S}_{6}$ ligands (see Supporting Information) in N-methylformamide (NMF) and 0.5 M $\mathrm{K}_{3}\mathrm{AsS}_{4}$ salt in NMF were prepared following a procedure previously outlined17. In a nitrogen filled glovebox, 400-500 $\mu$Ls of a solution of 5.8 nm PbS NCs with $\mathrm{K}_{4}\mathrm{Sn}_{2}\mathrm{S}_{6}$ ligands in mixtures of NMF and N,N-dimethylformamide (DMF) was loaded into a 2 mm path length quartz cuvette with custom 200 $\mu$m thick windows. A small stir bar was placed into the cuvette in the plane of the cuvette and the cuvette was then sealed using a rubber septum and parafilm. A syringe preloaded with a solution of 0.5 M $\mathrm{K}_{3}\mathrm{AsS}_{4}$ in NMF was attached to the cuvette via Teflon tubing through the septum. The tubing-septum interface was sealed with epoxy. The gas-tight apparatus was carefully moved into the beam path and the syringe placed onto a New Era syringe pump (model NE-1000). X-ray scattering data were collected continuously while stirring the solution using a magnetic stirrer from Ultrafast Systems. All X-ray scattering patterns were collected using $\sim$0.1--1 s exposures at a rate of one pattern every $\sim$0.1--5 s. For each in situ experiment, after about 5 min of data acquisition, the excess salt in NMF solution was injected using the syringe pump at a rate of $\sim$1 mL/s. The injection took $\sim$ 1 - 10 s depending on how much salt was added. The ionic strength of the solution post-quench is $I=\frac{1}{2}\sum_{n=1}^{N}c_{n}z_{n}^{2}$, where $c_{n}$ is the concentration of ion species $n$ in molar, $z_{n}$ is the valency of ion $n$, and $N$ is the number of different ion species in solution. The Debye length of the solution post-quench is $\lambda=\sqrt{\frac{\epsilon_{r}\epsilon_{0}k_{\mathrm{B}}T}{\sum_{n=1}^{N}c_{n}z_{n}^{2}e^{2}}}$ where $\epsilon_{r}$ is the solvent dielectric constant of 88, $\epsilon_{0}$ is the vacuum permittivity, and $e$ is the elemental charge. The apparatus was kept at room temperature (see Supporting Information for additional temperature considerations). Data were continuously acquired after injection for up to two hours. SAXS patterns of cuvettes filled with NMF/DMF and varying amounts of $\mathrm{K}_{3}\mathrm{AsS}_{4}$ salt were taken for background subtraction (Figure SLABEL:SI-background).
@@
Modeling of SAXS patterns. We model the background-subtracted scattered intensity as $I(q)=I_{\mathrm{colloid}}(q)+I_{\mathrm{liquid}}(q)+I_{\mathrm{SL}}(q)+I_{\mathrm{LQ}}(q)$, where $I_{\mathrm{colloid}}(q)$ is the scattered intensity from dilute, polydisperse, hard spheres, $I_{\mathrm{liquid}}(q)$ is the scattered intensity at high $q$ from a liquid of hard spheres, $I_{\mathrm{SL}}(q)$ is the scattered intensity at high $q$ from fcc SLs, and $I_{\mathrm{LQ}}(q)$ is the scattering at low $q$ due to the finite size of liquid and SL clusters. For $I_{\mathrm{colloid}}(q)$, $I_{\mathrm{SL}}(q)$, and $I_{\mathrm{LQ}}(q)$, we use models previously developed23. To compute $I_{\mathrm{liquid}}(q)$, we solve the Ornstein-Zernicke equation using the Percus-Yevick closure relation for the case of hard spheres55, 56. To explicitly compute $I_{\mathrm{colloid}}(q)$ and $I_{\mathrm{liquid}}(q)$ we use xrsdkit (https://github.com/scattering-central/xrsdkit). See supporting information for further details on the components of the model.
@@
Avrami analysis. We analyzed kinetics of liquid and SL formation using the Avrami equation, 58 $V(t)=1-\exp(kt^{n})$, where $V(t)$ is the volume fraction of the new phase. The Avrami exponent, $n$, which is typically between 1 and 4, describes the dimensionality of the system, with a contribution of 1 for nucleation and a contribution of 1 for each of the growth dimensions. We assume homogeneous nucleation and growth in 3 dimensions and thus expect an Avrami exponent of $n\sim 4$. For $n=4$, the Avrami rate constant, $k$, is $k=\frac{1}{3}k_{n}k_{g}^{3}$, where $k_{n}$ is the nucleation rate and $k_{g}$ is the growth rate in a single dimension. We use the fractions of NCs in the liquid and SL phases obtained from the scattering invariant calculation described above as proxies for $V(t)$, and fit these to the Avrami equation to extract $k$. Including density dependencies in the nucleating barrier and particle collision frequencies, classical nucleation theory predicts that $k$ depends on $\phi_{0}$ as $k=c\phi_{0}^{3}\exp(-\frac{a}{(\ln(\phi_{0}/\phi_{\mathrm{colloid}}))^{2}})$. 59, 60, 61 The parameters $a$, which depends on temperature and surface energy of the interface of the nucleus and the colloidal phase, $c$, which arises from the growth rate and the kinetic prefactor to the nucleation rate, and $\phi_{\mathrm{colloid}}$, which is the volume fraction of the colloidal phase at equilibrium, do not vary with $\phi_{0}$. See supporting information for further details.